# Fine-Tuning Multilingual Language Models for Code Review: An Empirical Study on Industrial C# Projects

### v1

@@ Abstract.
Code review is essential for maintaining software quality but often time-consuming and cognitively demanding, especially in industrial environments. Recent advancements in language models (LMs) have opened new avenues for automating core review tasks. This study presents the empirical evaluation of monolingual fine-tuning on the performance of open-source LMs across three key automated code review tasks: Code Change Quality Estimation, Review Comment Generation, and Code Refinement. We fine-tuned three distinct models—CodeReviewer, CodeLlama-7B, and DeepSeek-R1-Distill—on a C#-specific dataset combining public benchmarks with industrial repositories. Our study investigates how different configurations of programming languages and natural languages in the training data affect LM performance, particularly in comment generation. Additionally, we benchmark the fine-tuned models against an automated software analysis tool (ASAT) and human reviewers to evaluate their practical utility in real-world settings. Our results show that monolingual fine-tuning improves model accuracy and relevance compared to multilingual baselines. While LMs can effectively support code review workflows, especially for routine or repetitive tasks, human reviewers remain superior in handling semantically complex or context-sensitive changes. Our findings highlight the importance of language alignment and task-specific adaptation in optimizing LMs for automated code review.
@@ 1. Introduction
Further complicating this landscape, most prior evaluations have been conducted using open-source datasets drawn from platforms such as GitHub (Zhao et al., 2023; Li et al., 2022c). While these datasets offer scale and accessibility, they may not accurately reflect the complexity, coding standards, and domain-specific practices of industrial software development.
@@
To systematically explore model performance across these tasks, we pose the following research questions: RQ1. Can fine-tuning a multilingual code review–specific PLM (CodeReviewer) on monolingual C# data improve its performance on Code Change Quality Estimation and Code Refinement? RQ2. How does fine-tuning different types of LMs—CodeReviewer, CodeLlama, and DeepSeek—on different PL/NL combinations affect Review Comment Generation? RQ3. How do fine-tuned LMs compare to an ASAT and human reviewers in identifying review-worthy code changes and generating feedback?
@@
  • We conduct a systematic evaluation of three open-source LMs with distinct pretraining objectives (CodeReviewer, CodeLlama-7B, and DeepSeek-R1-Distill) fine-tuned on monolingual C# repositories across key ACR tasks.
  • We analyze the impact of dataset language composition (English-only vs. mixed-language vs. multilingual) on model performance, demonstrating how aligning training data with the target review language improves review comment generation quality.
  • We benchmark LMs against both an ASAT and human reviewers, providing a comprehensive assessment of their strengths and limitations across Code Change Quality Estimation, Review Comment Generation, and Code Refinement tasks.
  • We offer detailed, task-specific insights into model behavior in real-world industrial C# development settings.
@@ 3. Study Design and Methodology
[…]
Overview of the experimental workflow.
Figure 1. Overview of the experimental workflow.
@@ 3.2. Data Preparation
We built task-specific datasets by combining PRs and reviewer comments from five internal C# repositories at Lovion GmbH with samples from the CodeReviewer benchmark (Li et al., 2022a). Data extraction from Lovion’s repositories was conducted via Gitea’s REST API, ensuring comprehensive metadata coverage, including PR information, code diffs, and review comments.
To ensure data quality, we removed duplicate PR entries and filtered out incomplete or non-compilable code snippets. Since Lovion’s review comments were originally written in German, we translated them into English using machine translation to create a linguistically unified dataset compatible with the CodeReviewer benchmark. To ensure translation fidelity and domain relevance, we manually reviewed a stratified sample of the translated comments.
@@
For each of the three downstream tasks, we created dedicated datasets. Each was randomly split into training (85%), validation (7.5%), and test (7.5%) subsets. To mitigate ordering bias, all datasets were shuffled prior to splitting.
@@
Finally, we applied task-specific data formatting. For Code Change Quality Estimation, we labeled each diff hunk as either requiring a review comment ($y=1$) or not ($y=0$), ensuring a balanced class distribution. Positive samples ($y=1$) were directly sourced from the Comment Generation dataset. Negative samples ($y=0$) were extracted from the final PR versions where no reviewer comments were issued for the corresponding diff hunks, indicating either reviewer approval or irrelevance. For Comment Generation, each instance included a code diff and its corresponding human-written review comment. For Code Refinement, each data point consisted of a before-and-after code pair reflecting reviewer-suggested changes.
An overview of the C# data preparation pipeline—including extraction, cleaning, merging, and task-specific formatting—is provided in Figure 2. Detailed dataset statistics across all tasks and sources are summarized in Table 1.
[…]
Dataset preparation workflow.
Figure 2. Dataset preparation workflow.
@@ 3.4. Model Fine-Tuning
We fine-tuned three distinct models in this study: CodeReviewer (Li et al., 2022a), CodeLlama-7B (Roziere et al., 2023), and DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025). CodeReviewer is a transformer-based encoder-decoder PLM pretrained on multilingual code and review comment pairs. It was specifically developed for ACR tasks and thus represents a review-specialized PLM. CodeLlama-7B is a decoder-only LLM pretrained on a wide range of code corpora in multiple PLs. It serves as a code-pretrained LLM optimized for general code understanding and generation. DeepSeek-R1-Distill-Llama-8B is an instruction-tuned general-purpose LLM, distilled from LLaMA 3.1–8B, with no code-specific pretraining. It is designed to perform well across diverse NL tasks, including reasoning and instruction following.
This diversity in pretraining paradigms enables a comprehensive comparison of review-specific, code-oriented, and general-purpose LMs in realistic monolingual C# review settings.
Fine-Tuning CodeReviewer for Three ACR Tasks. We fine-tuned CodeReviewer for all three ACR tasks based on its original multi-task design and prior performance in both classification and generation tasks (Li et al., 2022a). We followed the original implementation as a basis and adapted task-specific hyperparameters where needed.
@@
    Review Comment Generation: We applied a two-stage fine-tuning approach. In the first stage, we used a mixed-language dataset that included German comments from the Lovion dataset and English comments from the CodeReviewer benchmark. In the second stage, we translated all German comments into English using Python’s googletrans library (Googletrans, [n. d.]) and retrained the model on a fully English dataset. Across both stages, the model was trained for approximately 3 epochs (around 7,500 steps) using a learning rate of $3\times 10^{-4}$ and a batch size of 6.
@@
Fine-Tuning CodeLlama and DeepSeek for Review Comment Generation. Given the linguistic complexity and practical relevance of the Review Comment Generation task, we fine-tuned both CodeLlama-7B and DeepSeek-R1-Distill-Llama-8B specifically for this purpose. Their instruction-tuned architectures and strong NL generation capabilities make them particularly well-suited for review comment generation. Due to computational constraints, we focused their evaluation exclusively on review comment generation to assess their ability to generate high-quality, context-aware review comments for C#.
@@
We fine-tuned both models using 4-bit QLoRA, a technique that enables efficient fine-tuning of large-scale models by combining weight quantization with low-rank adaptation. The shared hyperparameters included 3 training epochs, a LoRA rank of 32 (i.e., the dimensionality of the trainable low-rank matrices), a token limit of 2,048, a dropout rate of 0.05, and a learning rate of 0.0002, optimized using the paged AdamW algorithm. For CodeLlama-7B, we employed the Axolotl framework with a weight decay of 0.0, whereas for DeepSeek-R1-Distill-Llama-8B, we used the Unsloth framework with a weight decay of 0.01.
Training Data Scope and Language Composition. Since both the programming language (PL) used in the code and the natural language (NL) used in review comments vary across models and datasets, we classified the training configurations along two dimensions:
@@
All fine-tuning scripts, configuration files, and hyperparameter settings are publicly available in our replication package [^2] https://github.com/DAP5555/code-review.
@@ 3.5. SonarQube Integration
We used SonarQube as a rule-based baseline to evaluate the effectiveness of both LMs and human reviewers. Its outputs provided a reference point for identifying overlapping or missed issues across security, reliability, and maintainability dimensions. To integrate SonarQube into Lovion GmbH’s DevOps pipeline, we deployed the system within the company’s Jenkins-based continuous integration (CI) workflow (see Figure 3). This setup enabled automated analysis of all relevant PRs and ensured consistent quality assessment outputs. These outputs served as a static baseline against which we compared both LM-generated and human-written code review results.
[…]
Diagram showing SonarQube integration into the Jenkins CI pipeline.
Figure 3. SonarQube integration in the Jenkins CI pipeline.
@@ 3.6. Evaluation Framework
Evaluation Scope. To evaluate the effect of fine-tuning, we included both fine-tuned and non-fine-tuned versions of each model in our experimental setup. Our evaluation combines both task-wide assessments over the full dataset and human-aligned evaluations conducted on a 40-PR subset. While BLEU and execution time metrics were computed across all PRs, human-centered evaluations (e.g., Information, Relevance) and comparisons to human reviewers or SonarQube were restricted to the manually annotated subset. This dual-scope approach ensures both large-scale comparability and in-depth quality benchmarking.
Evaluation Metrics. We adopted a multi-metric framework to assess both task-specific model outputs and overall code review effectiveness.
@@
For Code Change Quality Estimation, we used standard classification metrics: accuracy, precision, recall, and F1 score.
For Review Comment Generation, we applied both:
  • Automated evaluation: We used BLEU-4 (Papineni et al., 2002) to measure lexical overlap between generated and reference comments. Following prior studies (Li et al., 2022a; Microsoft, 2020), we computed BLEU scores over the first 256 tokens of each output. For DeepSeek-R1-Distill-Llama-8B, we excluded chain-of-thought (CoT) reasoning segments prior to scoring to ensure comparability with other models.
@@
  • Human evaluation: Six professional software engineers rated each AI-generated comment along two dimensions: Information (extent to which the comment provides meaningful and actionable feedback) and Relevance (degree to which the comment addresses the actual issue in the code diff). Both dimensions were rated on a 5-point Likert scale (1-5). The final score for each comment was calculated by averaging the six individual ratings.
For Code Refinement, we used BLEU to assess n-gram similarity and EM to check for byte-level equality between generated and reference code revisions.

### v2

@@ Abstract.
Code review is essential for maintaining software quality but often time-consuming and cognitively demanding, especially in industrial environments. Recent advancements in language models (LMs) have opened new avenues for automating code review tasks. This study presents the empirical evaluation of monolingual fine-tuning on the performance of open-source LMs across three key automated code review tasks: Code Change Quality Estimation, Review Comment Generation, and Code Refinement. We fine-tuned three distinct models—CodeReviewer, CodeLlama-7B, and DeepSeek-R1-Distill—on a C#-specific dataset combining public benchmarks with industrial repositories. Our study investigates how different configurations of programming languages and natural languages in the training data affect LM performance, particularly in comment generation. Additionally, we benchmark the fine-tuned models against an automated static analysis tool (ASAT) and human reviewers to evaluate their practical utility in real-world settings. Our results show that monolingual fine-tuning improves model accuracy and relevance compared to multilingual baselines. While LMs can effectively support code review workflows, especially for routine or repetitive tasks, human reviewers remain superior in handling semantically complex or context-sensitive changes. Our findings highlight the importance of language alignment and task-specific adaptation in optimizing LMs for automated code review.
@@ 1. Introduction
Further complicating this landscape, most prior evaluations have relied on open-source datasets from platforms such as GitHub (Zhao et al., 2023; Li et al., 2022c). While these datasets offer scalability and ease of access, they often capture limited contextual diversity and may not accurately reflect the complexity, review workflows, and coding standards found in industrial software development environments.
@@
To systematically explore model performance across these tasks, we pose the following research questions: RQ1. How does monolingual fine-tuning affect the performance of LMs in detecting whether a code change requires human review? RQ2. How does fine-tuning different types of LMs on different PL/NL combinations affect review comment generation? RQ3. What is the impact of fine-tuning on the code refinement capability of LMs, in terms of producing accurate and functionally equivalent code revisions? RQ4. How do fine-tuned LMs compare to an ASAT and human reviewers in identifying review-worthy code changes and generating feedback?
@@
  • Cross-Paradigm Evaluation Framework: We systematically compare three open-source language models with distinct pretraining objectives—a review-specialized PLM (CodeReviewer), a code-pretrained LLM (CodeLlama-7B), and a general-purpose instruction-tuned LLM (DeepSeek-R1-Distill)—against both an industrial-strength ASAT (SonarQube) and human reviewers across all core ACR tasks.
  • Language-Aligned Fine-Tuning Protocol: We demonstrate that aligning both the programming language (monolingual C#) and natural language (English-only vs. bilingual vs. multilingual) of training data significantly impacts review comment quality.
  • Beyond Automated Metrics: We complement standard BLEU-based evaluation with expert-validated human assessments (Information, Relevance, Issue Correctness Rate) on 40 production PRs, revealing that lexical similarity often misaligns with practical review utility. This dual-scope evaluation—combining large-scale automated metrics with human-aligned quality measures—addresses known limitations of surface-level NLP metrics in code review contexts.
  • Industrial C# Benchmark: We introduce the first fine-tuned LM evaluation on enterprise-grade C# repositories, addressing a critical gap in ACR research which has predominantly focused on Java and Python in open-source settings.
@@ 3. Study Design and Methodology
Figure 1. Overview of the experimental workflow.:
@@ 3.2. Data Preparation
We built task-specific datasets by combining PRs and reviewer comments from five internal C# repositories at Lovion GmbH. Data extraction from Lovion’s repositories was conducted via Gitea’s REST API, ensuring comprehensive metadata coverage, including PR information, code diffs, and review comments.
These five repositories correspond to distinct modules of Lovion GmbH’s enterprise asset management platform, covering user interface, data access, visualization, workflow automation, and testing components. They differ in functionality and code complexity, providing a representative view of industrial C# development practices. All repositories are actively maintained under Lovion’s internal quality assurance policies, including mandatory peer review and continuous integration (CI) checks, ensuring that the extracted PRs and comments are high-quality and production-grade.
To complement the industrial dataset, we integrated the publicly available CodeReviewer benchmark (Li et al., 2022a). This addition enhanced data volume and supported stable model training. While the Lovion repositories provided rich, domain-specific review data representative of real industrial practices, their limited scale posed challenges for fine-tuning large transformer-based models, such as overfitting and unstable convergence. Incorporating the C# subset of the CodeReviewer benchmark mitigated these issues by supplying additional language-consistent samples without introducing domain overlap. To prevent data leakage, we included only C# examples and verified that both datasets originated from distinct platforms (Gitea vs. GitHub) with no shared PRs. All data were re-indexed, deduplicated, and filtered to remove incomplete or non-compilable code snippets before unification.
Translation and Quality Validation. Since the Lovion review comments were originally written in German, they were automatically translated into English using Python’s googletrans library (v4.0.0rc1, free API mode). To ensure translation fidelity, two bilingual annotators with professional English–German proficiency manually reviewed a stratified random sample of 200 translated comment pairs (6̃% of the dataset). Each pair was rated on a 1–5 adequacy and fluency scale, yielding mean scores of 4.6 and 4.8, respectively. Inter-annotator agreement was substantial (Cohen’s $\kappa$ = 0.82), indicating consistent judgments. Disagreements were resolved through discussion until consensus was reached. Minor domain-specific inconsistencies (e.g., technical terms and C#-specific vocabulary) were corrected during preprocessing, ensuring a high overall translation quality.
Language Configurations. The resulting dataset supports three NL configurations used across experiments: (i) English-only, consisting of translated Lovion comments and English samples from the CodeReviewer dataset; (ii) English+German, preserving both original and translated Lovion comments; and (iii) Multilingual, incorporating additional non-English samples (e.g., Chinese, Spanish, and French) from the CodeReviewer benchmark. Irrelevant or low-quality instances were removed during preprocessing to ensure consistency across configurations. These configurations are later used in the Review Comment Generation task to assess the effect of language composition on model performance.
@@
Data Splitting and Formatting. For each of the three downstream tasks, we created dedicated datasets. All PRs were first chronologically sorted by submission date. We then partitioned the data sequentially into training (85%), validation (7.5%), and test (7.5%) subsets, ensuring that newer PRs were never included in the training data relative to the test set. Within each subset, samples were randomly shuffled to mitigate potential ordering bias.
@@
Finally, we applied task-specific data formatting. For the Code Change Quality Estimation task, we constructed a total of 44,962 samples, equally divided between positive ($y=1$) and negative ($y=0$) instances. Positive samples ($y=1$) were directly sourced from the Comment Generation dataset, corresponding to diff hunks that received human review comments. Negative samples ($y=0$) were drawn from the final PR versions where no reviewer comments were issued, representing either accepted or non-critical changes. This balancing procedure helped mitigate bias toward the majority class and ensured comparable representation of both categories. For Comment Generation, each instance included a code diff and its corresponding human-written review comment. For Code Refinement, each data point consisted of a before-and-after code pair reflecting reviewer-suggested changes.
Detailed dataset statistics across all tasks and sources are summarized in Table 1.
@@ 3.4. Model Fine-tuning
We fine-tuned three distinct models in this study: CodeReviewer (Li et al., 2022a), CodeLlama-7B (Roziere et al., 2023), and DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025). CodeReviewer is a transformer-based encoder-decoder PLM pretrained on multilingual code and review comment pairs. It was specifically developed for ACR tasks and thus represents a review-specialized PLM. CodeLlama-7B is a decoder-only LLM pretrained on a wide range of code corpora in multiple PLs. It serves as a code-pretrained LLM optimized for general code understanding and generation. DeepSeek-R1-Distill-Llama-8B is an instruction-tuned general-purpose LLM, distilled from LLaMA 3.1–8B, with no code-specific pretraining. It is designed to perform well across diverse NL tasks, including reasoning and instruction following. This diversity in pretraining paradigms enables a comprehensive comparison of review-specific, code-oriented, and general-purpose LMs in monolingual C# review settings.
The selection of CodeReviewer, CodeLlama-7B, and DeepSeek-R1-Distill-Llama-8B was guided by three main criteria: (i) open-source accessibility ensuring full reproducibility and parameter-efficient fine-tuning; (ii) architectural diversity, covering a review-specific PLM (CodeReviewer), a code-pretrained LLM (CodeLlama), and a general-purpose instruction-tuned LLM (DeepSeek); and (iii) representativeness of SOTA open models widely adopted in automated code review research. Proprietary models such as GPT, Claude, or Gemini were not included due to closed-weight architectures, licensing restrictions, limited reproducibility in fine-tuning workflows and privacy concerns associated with uploading industrial code data to external APIs. Our focus is therefore on open, reproducible, and practically deployable models suitable for industrial integration.
3.4.1. Fine-Tuning CodeReviewer for Three ACR Tasks
We fine-tuned CodeReviewer for all three ACR tasks based on its original multi-task design and prior performance in both classification and generation tasks (Li et al., 2022a). We followed the original implementation as a basis and adapted task-specific hyperparameters where needed.
@@ 3.4.1. Fine-Tuning CodeReviewer for Three ACR Tasks
    Review Comment Generation: We applied a two-stage fine-tuning approach. In the first stage, we used a mixed-language dataset that included German comments from the Lovion dataset and English comments from the CodeReviewer benchmark. In the second stage, we used the translated version of the Lovion comments (see Section 3.2) to retrain the model on a fully English dataset. Across both stages, the model was trained for approximately 3 epochs (around 7,500 steps) using a learning rate of $3\times 10^{-4}$ and a batch size of 6.
@@
3.4.2. Fine-Tuning CodeLlama and DeepSeek for Review Comment Generation
In addition to the review-specialized CodeReviewer model, we fine-tuned both CodeLlama-7B and DeepSeek-R1-Distill-Llama-8B specifically for the Review Comment Generation task. Their instruction-tuned architectures and strong NL generation capabilities make them particularly well-suited for review comment generation. Due to computational constraints, we focused their evaluation exclusively on review comment generation to assess their ability to generate high-quality, context-aware review comments for C#.
@@ 3.4.2. Fine-Tuning CodeLlama and DeepSeek for Review Comment Generation
We fine-tuned CodeLlama and DeepSeek using 4-bit QLoRA, a technique that enables efficient fine-tuning of large-scale models by combining weight quantization with low-rank adaptation. The shared hyperparameters included 3 training epochs, a LoRA rank of 32 (i.e., the dimensionality of the trainable low-rank matrices), a token limit of 2,048, a dropout rate of 0.05, and a learning rate of 0.0002, optimized using the paged AdamW algorithm. For CodeLlama-7B, we employed the Axolotl framework with a weight decay of 0.0, whereas for DeepSeek-R1-Distill-Llama-8B, we used the Unsloth framework with a weight decay of 0.01.
3.4.3. Training Data Scope and Language Composition.
Since both the programming language (PL) used in the code and the natural language (NL) used in review comments vary across models and datasets, we classified the training configurations along two dimensions:
@@
3.5. Data and Code Availability
All scripts and fine-tuning configurations used in this study are available in the accompanying replication package. [^2] https://anonymous.4open.science/r/CodeReview-LMs/ The package contains preprocessing scripts, evaluation code, and instructions to reproduce all results reported in this paper. The industrial C# PR dataset from Lovion GmbH, however, cannot be publicly released due to confidentiality agreements with the company. To ensure transparency, we provide detailed metadata and aggregated statistics describing the dataset’s structure (e.g., issue types, PR size distribution, and module coverage) so that the experimental design and data characteristics can be replicated without accessing the raw code. Researchers interested in collaboration or controlled data access may contact the authors subject to company approval.
@@ 3.6. SonarQube Integration
While SonarQube was not designed to generate NL feedback, we included it as a complementary baseline to represent rule-based, non-linguistic analysis within our unified evaluation framework. This allowed us to systematically compare traditional static analysis with language-based review approaches in identifying review-relevant issues. To ensure comparability, we applied the same correctness criterion (i.e., “good” vs. “not good”) when evaluating its detections against ground truth. Hence, SonarQube serves not as a linguistic peer to LMs, but as a representative of established industrial practices for rule-based review support.
To integrate SonarQube into Lovion GmbH’s DevOps pipeline, we deployed the system within the company’s Jenkins-based continuous integration (CI) workflow. This setup enabled automated analysis of all relevant PRs and ensured consistent quality assessment outputs. These outputs served as a static baseline against which we compared both LM-generated and human-written code review results.
@@ 3.7. Evaluation Framework
To comprehensively address our research questions, we adopt a dual-scope evaluation: (i) a full test set for large-scale, fully automated metrics, and (ii) a 40-PR annotated subset for expert-validated, human-aligned analyses. This design provides both breadth (statistical comparability on all PRs) and depth (contextual quality against expert ground truth). We evaluate both fine-tuned and non–fine-tuned versions of each model.
3.7.1. Full test set (automated metrics)
PRs from real workflows with existing human comments were used to compute automated metrics across all PRs:
@@ 3.7.1. Full test set (automated metrics)
- For Code Change Quality Estimation, we used standard classification metrics—accuracy, precision, recall, and F1 score—to evaluate how accurately each model identified whether a PR required a review comment (problematic) or not (acceptable).
- For Review Comment Generation, we used BLEU-4 (Papineni et al., 2002) to measure lexical overlap between generated and reference comments. This choice aligns with the evaluation protocol in the CodeReviewer study (Li et al., 2022a), ensuring methodological comparability. Following prior works (Li et al., 2022a; Microsoft, 2020), BLEU-4 scores were computed over the first 256 output tokens to standardize sequence length and reduce the bias introduced by lengthy generations. For DeepSeek-R1-Distill-Llama-8B, chain-of-thought (CoT) reasoning segments were excluded prior to scoring for consistency. Although code-sensitive metrics such as CodeBLEU (Ren et al., 2020) or CrystalBLEU (Eghbali and Pradel, 2023) could offer complementary insights, they were not used because the original CodeReviewer checkpoints are not publicly available, making cross-study comparison infeasible.
- For Code Refinement, we used BLEU to assess n-gram similarity and Exact Match (EM) to verify byte-level equivalence between generated and reference code revisions. These metrics were selected to remain consistent with the CodeReviewer baseline setup and to enable one-to-one comparison of C#-specific results.
3.7.2. 40-PR annotated subset (human-aligned metrics)
We selected 40 PRs from Lovion GmbH’s industrial C# repositories using stratified random sampling to ensure balanced coverage across modules and issue types. Each PR contained a single code hunk to maintain a clear mapping between code change and review feedback. The subset consisted of 20 PRs that required review comments—covering issues categorized as bug, performance, maintainability, or security—and another 20 PRs that did not require any review comments, serving as a control group. Two senior engineers independently determined whether a review comment was required and assigned the appropriate issue category. Disagreements were resolved through discussion (Cohen’s $\kappa$ = 0.79). These expert-validated annotations established the ground truth foundation for all subsequent model and human performance evaluations.
Distinct evaluation protocols were applied for each ACR task:
- For Code Change Quality Estimation, LM and SonarQube predictions were compared against the expert-validated ground truth labels using standard classification metrics—accuracy, precision, recall, and F1 score. Human reviewers—six bilingual software engineers from Lovion GmbH (two senior, two mid-level, and two junior)—independently evaluated all 40 PRs to determine whether a review comment was required, and their majority decision served as the reference judgment for human performance.
- For Review Comment Generation, we used the 20 PRs that required comments. Two senior engineers (the same annotators who established the ground truth) evaluated whether each system’s output—either a natural-language comment (for humans and LMs) or a detected issue (for SonarQube)—correctly addressed the issue category defined in the ground truth. Each output was labeled as “correct” if it matched the issue type (e.g., bug, performance, maintainability, or security) and as “incorrect” otherwise. The proportion of correct outputs constituted the Issue Correctness Rate, providing a unified measure of issue-level correctness across rule-based and language-based systems.
@@
The six human reviewers also provided qualitative evaluations of AI-generated comments for the same 20 PRs. They rated each comment on two 5-point Likert scales: Information (extent to which the comment provides meaningful and actionable feedback) and Relevance (degree to which the comment addresses the actual issue in the code diff). The final score for each comment was calculated by averaging the six individual ratings. To avoid potential bias, the reviewers were independent of dataset curation and ground-truth annotation, and the two activities—manual code reviewing and evaluation of AI-generated outputs—were conducted separately to prevent cross-contamination of judgments. The aggregated human ratings thus represent a consensus-based human baseline that serves as the reference point across all evaluation dimensions.
3.7.3. Efficiency